OpenStudy (anonymous):

Find the domain of the following rational functions. y=(6x-5)/(x^2-6x+4) I got {XER} but I'm not sure if that's correct.

OpenStudy (anonymous):

by domain you just mean whether it's all real numbers or not correct?

OpenStudy (anonymous):

Yes, basically whether there are any limitations on the x-values.

hartnn (hartnn):

note that denominator cannot be = 0 so, exclude those values for which (x^2-6x+4) =0

OpenStudy (anonymous):

I figured that since it cannot be factored, it has no solution, and therefore results in all real numbers. Unless there IS a way to solve it?

OpenStudy (anonymous):

even if you use the quadratic formula?

hartnn (hartnn):

there is a solution, but not rational, irrational roots are there

hartnn (hartnn):

and you need to exclude those from real domain too

OpenStudy (anonymous):

[6+/-\sqrt{20}/2 \]

OpenStudy (anonymous):

not sure if the answer would be left in radical form

hartnn (hartnn):

= \(3\pm \sqrt 5\)

hartnn (hartnn):

so, you domain will be all real values of x , excluding x = 3 -sqrt 5 , x = 3 +sqrt 5

OpenStudy (anonymous):

But how did it go from \[(6\pm \sqrt{20})\div2\] to 3 +/-sqrt 5

hartnn (hartnn):

\(\Large \sqrt{20}=\sqrt {5\times4 }=\sqrt 4\sqrt 5=2\sqrt5\)

hartnn (hartnn):

then cancelling out 2 from numerator and denominator

OpenStudy (anonymous):

Ohh, right. Thank you! I remember doing that last year, but we haven't reviewed it again this year.

hartnn (hartnn):

welcome ^_^